# d2Cache: Accelerating Diffusion-Based LLMs via Dual Adaptive Caching

### v1

@@ 1 Introduction
Diffusion models have recently achieved remarkable success in generating continuous data like images (Yang et al., 2023), but text generation—a fundamentally discrete task—has long been dominated by autoregressive models (ARMs) (Touvron et al., 2023; Achiam et al., 2023; Guo et al., 2025). Building on the foundations of ARMs, recent studies have successfully extended diffusion processes to discrete language modeling and further scaled up these models (Nie et al., 2025; Ye et al., 2025; Li et al., 2025). These diffusion-based large language models (dLLMs) offer several key advantages over ARMs, such as mitigating the “reversal curse” (Berglund et al., 2023) and capturing high-level global semantic patterns (Nagarajan et al., 2025).
@@
To address the above efficiency challenges, recent studies (Ma et al., 2025; Wu et al., 2025; Liu et al., 2025; Hu et al., 2025) have explored approximate KV cache to accelerate dLLM inference. These studies build on the following key observation: for a subset of tokens, their KV states often exhibit high similarity across consecutive decoding steps. This enables to approximately reuse these KV states, which can avoid redundant computations and reduce the overall inference cost. In practice, these studies typically divide the sequence (including prompt tokens, masked tokens, and decoded tokens) into a static segment, where their KV states can be approximately reused, and a dynamic segment, where their KV states need to be frequently updated within a fixed window of decoding steps. However, these studies are coarse-grained and apply the same strategy to all tokens within both static and dynamic segments. As a result, they either suffer from limited flexibility or require complicated tuning. Moreover, since coarse-grained designs cannot capture the fine-grained token-level dynamics of KV states, they inevitably reuse KV states that should be updated, or update KV states that can be safely reused, thus limiting the achievable acceleration gains.
@@
To address these limitations, we seek to develop an effective fine-grained approximate KV cache strategy, which can adaptively select tokens and update their KV states at each decoding step rather than within a fixed decoding window (Ma et al., 2025; Wu et al., 2025; Liu et al., 2025; Hu et al., 2025). To this end, we first perform a fine-grained analysis to investigate the KV state dynamics in dLLMs. Our results show that, for masked tokens, their KV states evolve through three phases: (1) a gradual-change phase during the early decoding steps, (2) a rapid-change phase in the few steps immediately preceding their decoding, and (3) a stable phase after being decoded. Notably, we find that it is sufficient to update the KV states of masked tokens only during the rapid-change phase.
@@ 2 Related work
Diffusion-based large language models. Building on the success of diffusion models in continuous domains, such as image and video generation (Yang et al., 2023; Ho et al., 2022), recent studies have extended diffusion models to discrete language tasks (Sahoo et al., 2024; Shi et al., 2024; Nie et al., 2024; Arriola et al., 2025). Unlike autoregressive models (ARMs) that generate tokens sequentially (Touvron et al., 2023; Achiam et al., 2023; Guo et al., 2025), dLLMs feature an iterative denoising process over masked sequences, which can enable bidirectional context modeling and inherently support parallel decoding (Li et al., 2025). More recently, large-scale dLLMs, such as LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025), have demonstrated competitive performance on reasoning and instruction-following tasks, establishing themselves as a promising alternative to ARMs. Despite their promising performance, their reliance on bidirectional attention necessitates substantial inference overheads, which significantly hinder their practical deployments.
Approximate KV cache for dLLMs. Due to bidirectional attention, dLLMs cannot directly benefit from the standard KV cache (Li et al., 2025) as ARMs do. To address this limitation, recent studies have observed that the KV states in dLLMs remain highly similar across consecutive decoding steps. Building on this observation, several approximate KV caching techniques have recently emerged (Liu et al., 2025; Ma et al., 2025; Wu et al., 2025; Hu et al., 2025). Among them, dLLM-Cache (Liu et al., 2025) partitions the input sequence into two segments—prompt and response—and updates their KV states at different frequencies. dKV-Cache (Ma et al., 2025) introduces a one-step delayed KV caching scheme, in which decoded tokens are stored not at the current decoding step but at the subsequent decoding step. Fast-dLLM (Wu et al., 2025) features block-wise semi-autoregressive decoding and caches all KV states except those in the current decoding block. However, due to the coarse-grained nature, these methods inevitably reuse KV states that should be actively updated or update KV states that can be safely reused, which thus suffer from inferior acceleration gains. A comprehensive comparison between our d2Cache and two concurrent similar works (i.e., dLLM-Cache and Fast-dLLM) is provided in appendix A of the Appendix.
@@ 3.1 Generation process of dLLMs
dLLMs feature an iterative denoising paradigm to generate text over $T$ discrete decoding steps (Nie et al., 2025), where a fully masked initial sequence is progressively transformed into a fully unmasked final output. Formally, let $\mathcal{V}$ denote the token vocabulary, which includes a special token [MASK]. The inference process begins with an initial sequence $y_{0}$ of length $L$, which is constructed by concatenating a prompt segment $p$ with a response segment $r_{0}$ of $n$ masked tokens. We denote the set of indices corresponding to these masked tokens as $M_{0}=\{\lvert p\rvert,\lvert p\rvert+1,\dots,\lvert p\rvert+n-1\}$.
At each decoding step $t\in[0,\ldots,T-1]$, the corresponding sequence $y_{t}$ is first fed into the given dLLM model as input, which produces a probability distribution $p(x_{t}^{i}\mid y_{t})$ over the vocabulary for each masked position $x_{t}^{i}$. Based on this distribution, the most confident token predictions $\hat{X}_{t}$ and their associated confidence scores $S_{t}$ can be derived as follows:
@@
In practice, the scheduling is typically performed either by randomly sampling a subset of $M_{t}$ or by choosing those masked positions with the highest confidence scores (Nie et al., 2025). Subsequently, the masked index set for the next decoding step is updated as $M_{t+1}=M_{t}\setminus I_{t}$. After $T$ iterations, when the condition $M_{T}=\emptyset$ holds, the whole generation process is stopped and we get the final sequence $y_{T}$ with no remaining masked tokens (Nie et al., 2025).
@@
Figure 2: (a) PCA of 77th masked token’s trajectory on LLaMA-8B-Instruct with GSM8K (LL=328, nn=256, and TT=256). (b) Sequential distances between token pairs decoded in adjacent steps.
@@ 3.2 KV state dynamics and decoding order in dLLMs
Decoding order in dLLMs. Building on the above findings, a natural question arises: how can we determine whether a masked token is about to be decoded before its actual decoding—essentially a “chicken-and-egg” problem? To shed light on this, we leverage LLaDA-8B-Instruct and randomly sample 64 examples from GSM8K, in which we analyze the sequential distance between token pairs decoded in adjacent steps. As shown in Figure 2 (b), LLaDA-8B-Instruct tends to decode the next masked token from positions close to the most recently decoded token, with 90% of tokens falling within a distance of 10. This reveals an interesting decoding pattern: dLLMs tend to decode masked tokens located near previously decoded tokens. Therefore, we can estimate whether a masked token is about to be decoded according to the density of decoded tokens in its local context.
@@ 3.3 Attention distributions in dLLMs
Prior research on ARMs has observed that attention is not uniformly distributed but instead concentrated on a small subset of salient tokens (Xiao et al., 2023). This observation has served as the foundation for various optimization techniques, which apply differentiated strategies to tokens based on their importance (Feng et al., 2024; Cai et al., 2024). This naturally raises the following question: can the above observation from ARMs generalize to dLLMs? To answer this question, we conduct experiments on LLaDA-8B-Instruct with GSM8K to analyze the attention distribution.
Attention salience among tokens. Inspired by prior attention studies on ARMs, we employ attention rollout (Abnar & Zuidema, 2020) to visualize how attention propagates across tokens. The attention rollout algorithm aggregates cumulative attention by recursively multiplying the attention matrices across layers, yielding a global attribution map that highlights how information propagates from input tokens to the final output. More details about the attention rollout algorithm are provided in Section 4.2. As shown in Figure 3 (a, bottom), queries consistently attend to a small subset of key positions in prompt and decoded tokens, revealing that these tokens dominate the attention distribution compared to other tokens. As shown in Figure 3 (a, top), masked tokens receive negligible attention, which is substantially lower than that allocated to both prompt and decoded tokens.
@@
Key Observations and Takeaways 1. The KV states of masked tokens evolve through three phases: gradual change $\rightarrow$ rapid change $\rightarrow$ stable. Notably, it is sufficient to update the KV states of masked tokens only during the rapid-change phase, whereas the KV states of masked tokens can be safely cached for reuse during the gradual-change and stable phases.   Figure 2 (a) 2. The decoding order of masked tokens correlates with the number of decoded tokens in the local context. Therefore, we can estimate whether the masked token is in the rapid-change phase according to the density of decoded tokens in the local context.   Figure 2 (b) 3. The attention distribution in dLLMs is highly uneven: prompt and decoded tokens receive most attention from other tokens, while masked tokens only receive negligible attention. Moreover, attention allocations across consecutive decoding steps are highly similar. The attention allocation of the current step can thus be used to select tokens for KV state updates in the next step, i.e., prioritizing tokens that receive greater attention.   Figure 3
@@
Motivated by the observations in Section 3, we present Dual aDaptive Cache (d2Cache), a training-free approximate KV cache framework for accelerating dLLM inference. Unlike ARMs, which can naturally reuse previous KV states (Li et al., 2024), dLLMs cannot exploit this mechanism due to their non-autoregressive decoding nature (Wu et al., 2025), as shown in Figure 1. To bridge this gap, d2Cache seeks to adaptively identify tokens whose KV states should be actively updated at each decoding step, while caching the remaining tokens for reuse in subsequent decoding step.
Overview of d2Cache. As seen in (Nie et al., 2025), tokens in dLLMs can be grouped into three categories: prompt tokens, masked tokens, and decoded tokens. Based on this categorization, we introduce a two-stage fine-grained token selection strategy. ❶ Certainty prior-guided selection from masked tokens. After each forward pass, d2Cache assigns each masked token a certainty prior, defined as the product of its prediction confidence and the density of known tokens (i.e., prompt or decoded tokens) in its local context. d2Cache then adaptively selects a subset of masked tokens with higher certainty prior. In light of this, d2Cache naturally delivers an alternative decoding scheme: masked tokens can be decoded according to their certainty prior rather than prediction confidence. This certainty prior-guided decoding has proven more reliable than the default confidence-based decoding (see Table 2). ❷ Attention-aware selection from remaining tokens. Furthermore, for the remaining tokens (especially prompt and decoded tokens), d2Cache adaptively selects a subset of tokens with higher attention activations, which can be identified using attention rollout Abnar & Zuidema (2020). Finally, for the tokens selected in these two stages, d2Cache updates their KV states at each decoding step, while caching the KV states of the remaining tokens for reuse in subsequent decoding step. An intuitive example of this two-stage token selection is provided in Figure 1 (c).
@@ 4.1 Stage 1: certainty prior-guided selection
As shown in Figure 2 (b), the decoding order in dLLMs is highly localized: 90% of subsequent tokens are decoded within a distance of 10 from the most recently decoded token. Building on this finding, we introduce certainty prior, which quantifies (1) the prediction confidence and (2) the certainty density of neighboring tokens that are known (i.e., prompt or decoded tokens). For each masked token, we define its certainty prior as the product of its prediction confidence and the density of known tokens in its local context. In practice, the certainty prior can capture structural certainty, where higher value indicates that the masked token is more likely to be decoded sooner.
@@ 4.2 Stage 2: attention-aware selection
In Section 4.1, we present certainty prior-guided selection, which explores masked tokens whose KV states should be updated at each decoding step. In this section, we extend the selection process to the remaining tokens. Notably, we observe that attention rollout (Abnar & Zuidema, 2020)—a widely used attention analysis technique in ARMs—can effectively generalize to dLLMs, particularly for analyzing prompt and decoded tokens, making it well suited for our subsequent token selection.
As described in (Abnar & Zuidema, 2020), the attention rollout algorithm aggregates cumulative attention by recursively multiplying the attention matrices across layers, yielding a global distribution map that reveals how information propagates from input tokens to the final output. Formally, let $U$ denote the indices of the remaining tokens. At the decoding step $t+1$, the input of the given dLLM is no longer the full sequence $y_{t+1}$, but instead a subset of it:
@@
where $\bm{e}_{i}$ is the one-hot vector with a value of 1 at position $i$. Following (Abnar & Zuidema, 2020), we further define the per-layer transition matrix $W^{(l)}$ by combining the expanded attention matrix $E^{(l)}$ with the residual connection (i.e., an identity matrix $I$) and applying row-wise normalization:
@@ 5 Experiments
MBPP
3-shot
Gen. Len. = 512
Vanilla 2.48 (1.0×) 199.90 14.4 2.73 (1.0×) 182.78 52.0
+ dLLM-Cache 6.97 (2.8×) 71.79 12.8 7.07 (2.6×) 71.13 52.4
+ Fast-dLLM 6.80 (2.7×) 73.27 13.8 7.29 (2.7×) 69.47 52.0
@@
HumanEval
0-shot
Gen. Len. = 512
Vanilla 4.99 (1.0×) 105.76 45.1 4.39 (1.0×) 114.86 56.7
@@
Math-500
4-shot
Gen. Len. = 256
Vanilla 3.08 (1.0×) 82.51 38.4 3.51 (1.0×) 71.05 45.2
@@
AVG
Vanilla 3.33 (1.0×) 124.61 43.9 3.31 (1.0×) 113.66 57.6
+ dLLM-Cache 7.66 (2.3×) 49.36 43.1 6.78 (2.0×) 58.64 56.9
+ Fast-dLLM 8.74 (2.6×) 46.58 43.2 9.01 (2.7×) 45.43 57.4

### v2

@@ 1 Introduction
Diffusion models have recently achieved remarkable success in generating continuous data like images (Yang et al., 2023), but text generation—a fundamentally discrete task—has long been dominated by autoregressive models (ARMs) (Touvron et al., 2023; Achiam et al., 2023; Guo et al., 2025). Building on the foundations of ARMs, recent studies have successfully extended diffusion processes to discrete language modeling and further scaled up these models (Nie et al., 2025; Ye et al., 2025; Li et al., 2025c). These diffusion-based large language models (dLLMs) offer several key advantages over ARMs, such as mitigating the “reversal curse” (Berglund et al., 2023) and capturing high-level global semantic patterns (Nagarajan et al., 2025).
@@
To address the above efficiency challenges, recent studies (Ma et al., 2025; Wu et al., 2025; Liu et al., 2025; Hu et al., 2025) have explored approximate KV cache to accelerate dLLM inference. These studies build on the following key observation: for a subset of tokens, their KV states often exhibit high similarity across consecutive decoding steps. This enables to approximately reuse these KV states, which can avoid redundant computations and reduce the overall inference cost. In practice, these studies typically divide the sequence into a static segment, where their KV states can be approximately reused, and a dynamic segment, where their KV states need to be frequently updated within a fixed window of decoding steps. However, these studies are coarse-grained and apply the same strategy to all tokens within both static and dynamic segments. As a result, they either suffer from limited flexibility (Wu et al., 2025) or require complicated tuning (Liu et al., 2025). Moreover, since coarse-grained designs cannot capture the fine-grained token-level dynamics of KV states, they inevitably reuse KV states that should be updated, or update KV states that can be safely reused, thus limiting the achievable acceleration gains.
@@
To address these limitations, we seek to develop an effective fine-grained approximate KV cache strategy, which can adaptively select tokens and update their KV states at each decoding step rather than within a fixed decoding window. To this end, we first perform a fine-grained analysis to investigate the KV state dynamics in dLLMs. Our results show that, for masked tokens, their KV states evolve through three phases: (1) a gradual-change phase during the early decoding steps, (2) a rapid-change phase in the few steps immediately preceding their decoding, and (3) a stable phase after being decoded. Notably, we find that it is sufficient to update the KV states of masked tokens only during the rapid-change phase.
@@ 2 Related work
Diffusion-based large language models (dLLMs). Building on the success of diffusion models in continuous domains, such as image and video generation (Yang et al., 2023; Ho et al., 2022), recent studies have extended diffusion models to discrete language tasks (Sahoo et al., 2024; Shi et al., 2024; Nie et al., 2024; Arriola et al., 2025). Unlike autoregressive models (ARMs) that generate tokens sequentially (Touvron et al., 2023; Achiam et al., 2023; Guo et al., 2025), dLLMs feature an iterative denoising process over masked sequences, which can enable bidirectional context modeling and inherently support parallel decoding (Li et al., 2025c; a; b). More recently, large-scale dLLMs, such as LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025), have demonstrated competitive performance on reasoning and instruction-following tasks, establishing themselves as a promising alternative to ARMs. Despite their promising performance, their reliance on bidirectional attention necessitates substantial inference overheads, which significantly hinder their practical deployments.
Approximate KV cache for dLLMs. Due to bidirectional attention, dLLMs cannot directly benefit from the standard KV cache (Li et al., 2025c) as ARMs do. To address this limitation, recent studies have observed that the KV states in dLLMs remain highly similar across consecutive decoding steps. Building on this observation, several approximate KV caching techniques have recently emerged (Liu et al., 2025; Ma et al., 2025; Wu et al., 2025; Hu et al., 2025). Among them, dLLM-Cache (Liu et al., 2025) partitions the input sequence into two segments—prompt and response—and updates their KV states at different frequencies. dKV-Cache (Ma et al., 2025) introduces a one-step delayed KV caching scheme, in which decoded tokens are stored not at the current decoding step but at the subsequent decoding step. Fast-dLLM (Wu et al., 2025) features block-wise semi-autoregressive decoding and caches all KV states except those in the current decoding block. However, due to the coarse-grained nature, these methods inevitably reuse KV states that should be actively updated or update KV states that can be safely reused, which thus suffer from inferior acceleration gains. A comprehensive comparison between our d2Cache and two concurrent similar works (i.e., dLLM-Cache and Fast-dLLM) is provided in appendix B of the Appendix.
@@ 3.1 Generation process of dLLMs
As shown in Nie et al. (2025), dLLMs feature an iterative denoising paradigm to generate text over $T$ discrete decoding steps, where a fully masked initial sequence is progressively transformed into a fully unmasked final output. Formally, let $\mathcal{V}$ denote the token vocabulary, which includes a special masked token [MASK]. The inference process of dLLMs begins with an initial sequence $y_{0}$ of length $L$, which is simply constructed by concatenating a prompt segment $p$ with a response segment $r_{0}$ that consists of $n$ masked tokens. We denote the set of indices corresponding to these masked tokens as $M_{0}=\{\lvert p\rvert,\lvert p\rvert+1,\dots,\lvert p\rvert+n-1\}$.
At each decoding step $t\in[0,\ldots,T-1]$, the corresponding sequence $y_{t}$ is first fed into the given dLLM as input, which produces a probability distribution $p(x_{t}^{i}\mid y_{t})$ over the vocabulary for each masked position $x_{t}^{i}$. Based on this distribution, the most confident token predictions $\hat{X}_{t}$ and their associated confidence scores $S_{t}$ can be derived as follows:
@@
In practice, the scheduling is typically performed either by randomly sampling a subset of $M_{t}$ or by choosing those masked positions with the highest confidence scores (Nie et al., 2025). Subsequently, the masked index set for the next decoding step is updated as $M_{t+1}=M_{t}\setminus I_{t}$. After $T$ iterations, when the condition $M_{T}=\emptyset$ holds, the whole generation process is stopped and we get the final sequence $y_{T}$ with no remaining masked tokens.
@@
Figure 2: (a) PCA visualization of 77th masked token’s trajectory on LLaDA-8B-Instruct with GSM8K (LL=328, nn=256, and TT=256). (b) Sequential distances between token pairs decoded in adjacent steps.
@@ 3.2 KV state dynamics and decoding order in dLLMs
Decoding order in dLLMs. Building on the above findings, a natural question arises: how can we determine whether a masked token is about to be decoded before its actual decoding—essentially a “chicken-and-egg” problem? To shed light on this, we randomly sample 64 examples from GSM8K, in which we analyze the sequential distance between token pairs decoded in adjacent steps. As shown in Figure 2 (b), LLaDA-8B-Instruct tends to decode the next masked token from positions close to the most recently decoded token, with 90% of tokens falling within a distance of 10. This reveals an interesting decoding pattern: dLLMs tend to decode masked tokens located near previously decoded tokens. Therefore, we can estimate whether a masked token is about to be decoded according to the density of decoded tokens in its local context.
@@ 3.3 Attention distributions in dLLMs
Prior research on ARMs has observed that attention is not uniformly distributed but instead concentrated on a small subset of salient tokens (Xiao et al., 2023). This observation has served as the foundation for various optimization techniques, which apply differentiated strategies to tokens based on their importance (Feng et al., 2024; Cai et al., 2024; Zhu et al., 2024). This naturally raises the following question: can the above observation from ARMs generalize to dLLMs? To answer this question, we conduct experiments on LLaDA-8B-Instruct with GSM8K to analyze the attention distribution.
Attention salience among tokens. Inspired by prior attention studies on ARMs, we employ attention rollout (Abnar and Zuidema, 2020) to visualize how attention propagates across tokens. The attention rollout algorithm aggregates cumulative attention by recursively multiplying the attention matrices across layers, yielding a global attribution map that highlights how information propagates from input tokens to the final output. More details about the attention rollout algorithm are provided in Section 4.2. As shown in Figure 3 (a, bottom), queries consistently attend to a small subset of key positions in prompt and decoded tokens, revealing that these tokens dominate the attention distribution compared to other tokens. As shown in Figure 3 (a, top), masked tokens receive negligible attention, which is substantially lower than that allocated to both prompt and decoded tokens.
@@
Motivated by the observations in Section 3, we present Dual aDaptive Cache (d2Cache), a training-free approximate KV cache framework for accelerating dLLM inference. To exploit approximate KV cache in dLLMs, d2Cache seeks to adaptively identify tokens whose KV states should be actively updated at each decoding step, while caching the remaining tokens for reuse in a subsequent decoding step.
Overview of d2Cache. We first group tokens in dLLMs into three categories: prompt tokens, masked tokens, and decoded tokens. Based on this categorization, we introduce a two-stage fine-grained token selection strategy. ❶ Certainty prior-guided selection from masked tokens. After each forward pass, d2Cache assigns each masked token a score combining prediction confidence and a certainty prior, which represents the density of known tokens (i.e., prompt or decoded ones) in the local context. d2Cache then adaptively selects a subset of masked tokens with higher scores. In light of this, d2Cache naturally enables an alternative decoding scheme guided by the certainty prior rather than prediction confidence alone, yielding higher reliability (see Table 2). ❷ Attention-aware selection from remaining tokens. Furthermore, for the remaining tokens (especially prompt and decoded tokens), d2Cache adaptively selects a subset of tokens with higher attention activations. Finally, for the tokens selected in these two stages, d2Cache updates their KV states at each decoding step, while caching the KV states of the remaining tokens for reuse in a subsequent decoding step. An intuitive example of this two-stage token selection is provided in Figure 1 (c).
@@ 4.1 Stage 1: certainty prior-guided selection
As shown in Figure 2 (b), the decoding order in dLLMs is highly localized: 90% of subsequent tokens are decoded within a distance of 10 from the most recently decoded token. Building on this finding, we introduce a certainty prior, defined as the density of known tokens (i.e., prompt or decoded tokens) in the local context. In practice, the certainty prior can capture structural certainty, where higher value indicates that the masked token is more likely to be decoded sooner.
@@ 4.2 Stage 2: attention-aware selection
In Section 4.1, we present certainty prior-guided selection, which explores masked tokens whose KV states should be updated at each decoding step. In this section, we extend the selection process to the remaining tokens. Notably, we observe that attention rollout (Abnar and Zuidema, 2020)—a widely used attention analysis technique in ARMs—can effectively generalize to dLLMs, particularly for analyzing prompt and decoded tokens, making it well suited for our subsequent token selection.
As described in Abnar and Zuidema (2020), the attention rollout algorithm aggregates cumulative attention by recursively multiplying the attention matrices across layers, yielding a global distribution map that reveals how information propagates from input tokens to the final output. Formally, let $U$ denote the indices of the remaining tokens. At the decoding step $t+1$, the input of the given dLLM is no longer the full sequence $y_{t+1}$, but instead a subset of it:
@@
where $\bm{e}_{i}$ is the one-hot vector with a value of 1 at position $i$. Following Abnar and Zuidema (2020), we further define the per-layer transition matrix $W^{(l)}$ by combining the expanded attention matrix $E^{(l)}$ with the residual connection (i.e., an identity matrix $I$) and applying row-wise normalization:
@@ 5 Experiments
MBPP 3-shot Gen. Len. = 512 Vanilla 2.48 (1.0×) 199.90 38.0 2.73 (1.0×) 182.78 52.0
+ dLLM-Cache 6.97 (2.8×) 71.79 38.0 7.07 (2.6×) 71.13 52.4
+ Fast-dLLM 6.80 (2.7×) 73.27 38.4 7.29 (2.7×) 69.47 52.0
@@
HumanEval 0-shot Gen. Len. = 512 Vanilla 4.99 (1.0×) 105.76 45.1 4.39 (1.0×) 114.86 56.7
@@
Math-500 4-shot Gen. Len. = 256 Vanilla 3.08 (1.0×) 82.51 38.4 3.51 (1.0×) 71.05 45.2
@@
GPQA 0-shot Gen. Len. = 256 Vanilla 6.14 (1.0×) 43.34 25.2 6.43 (1.0×) 41.14 30.1
+ dLLM-Cache 11.51 (1.9×) 22.33 27.2 10.91 (1.7×) 23.62 31.0
+ Fast-dLLM 12.41 (2.0×) 20.66 25.7 11.75 (1.8×) 21.79 34.6
[…]
MMLU-Pro 5-shot Gen. Len. = 256 Vanilla 1.76 (1.0×) 152.62 37.5 2.15 (1.0×) 126.31 47.9
+ dLLM-Cache 6.79 (3.9×) 38.29 38.1 7.82 (3.6×) 34.09 46.5
+ Fast-dLLM 8.91 (5.1×) 29.00 37.1 9.74 (4.5×) 27.69 45.9
[…]
AVG Vanilla 3.54 (1.0×) 115.73 43.6 3.64 (1.0×) 103.68 51.4
+ dLLM-Cache 8.16 (2.3×) 43.01 43.8 7.64 (2.1×) 48.71 50.9
+ Fast-dLLM 9.38 (2.7×) 39.33 43.4 9.59 (2.6×) 38.53 51.7